OpenStudy (anonymous):

-4(-7w + 6) -9w=4(w-6) -9 solve for W

OpenStudy (anonymous):

-4*-7=28 since a negative number * a negative number = a positive number and -4*6= -24 since a negative number*a positive number = a negative number so 28w-24-9w=4w-24-9 so move the 4w to the LHS and the -24 to the RHS to get: 28w-9w-4w=-24-9+24 so now just collect the terms and find w